# Observations of high-frequency spectral peaks from in-situ waves in ice data: evidence for nonlinear waves in ice triad interactions?

### v1

@@ 5 Conclusion
6 Acknowledgments
The help of the UNIS logistics team during the 2015 T15 fieldwork is gratefully acknowledged. The help of laboratory engineer Olav Gundersen (UiO) when building and preparing the equipment used in the T15 and Y20 deployments is gratefully acknowledged. We would like to thank the crew of the RV Kronprins Haakon for invaluable help deploying the buoys in the Arctic. The authors thank Pr. Yngve Kristoffersen (UiB) for having deployed the buoy used in the Y20 dataset. Funding from the Research Council of Norway through the projects "Waves in Oil and Ice" (WOICE, project 233901), "Dynamics Of Floating Ice" (DOFI, project 280625), "Arven Etter Nansen" (AeN, project 276730), "FOCUS" (project 301450), is gratefully acknowledged.

### v5

@@ 6 Use of Artificial Intelligence tools in the writing of this manuscript
[…]
Conceptualization: J.R. Data collection and curation: J.R., J.V., T.N., G.S., T.K., A.J., Ø.B., G.H., M.M., A.M., B.W., K.H.C., P.H. Formal analysis and Software: J.R., J.V. Funding acquisition: J.R., J.V., A.B., T.W., A.J., Ø.B., M.M., K.H.C. Investigation: J.R., J.V., T.N., G.S., A.B., T.W., L.W.D., A.M., K.T. Methodology: J.R, A.B., T.W., J.V, T.N. Project administration: J.R., A.B., T.W., A.J., Ø.B., M.M., L.W., A.M., P.H. Validation: J.R., J.V., T.N. Visualization: J.R., J.V., T.N. Writing - original draft preparation: J.R., with contributions from from J.V. Writing - review & editing: all.
[…]
Some authors are members of the editorial board of the journal "The Cryosphere".
Acknowledgements.
The help of the UNIS logistics team during the 2015 T15 fieldwork is gratefully acknowledged. The help of laboratory engineer Olav Gundersen (UiO) when building and preparing the equipment used in the T15 and Y20 deployments is gratefully acknowledged. We would like to thank the crew of the RV Kronprins Haakon for invaluable help deploying the buoys in the Arctic. The authors thank Pr. Yngve Kristoffersen (UiB) for having deployed the buoy used in the Y20 dataset. Funding from the Research Council of Norway through the projects "Waves in Oil and Ice" (WOICE, project 233901), "Dynamics Of Floating Ice" (DOFI, project 280625), "Arven Etter Nansen" (AeN, project 276730), "FOCUS" (project 301450), is gratefully acknowledged. Support was also received from the Australian Antarctic Program (AAS4496, AAS4506, AAS4625) and the Australian government (ASCI000002).